OpenStudy (anonymous):

If f(x)=x-5, g(x)=-3x, and h(x)=2x^2-6x+1, find the following. h(-1)

OpenStudy (dumbcow):

substitute -1 for x -> 2(-1)^2 -6(-1) +1 =2+6+1 = 9